# IMAGHarmony: Controllable Image Editing with Consistent Object Quantity and Layout

### v1

@@ Abstract
Recent diffusion models have advanced image editing by enhancing visual quality and control, supporting broad applications across creative and personalized domains. However, current image editing largely overlooks multi-object scenarios, where precise control over object categories, counts, and spatial layouts remains a significant challenge. To address this, we introduce a new task, quantity-and-layout consistent image editing (QL-Edit), which aims to enable fine-grained control of object quantity and spatial structure in complex scenes. We further propose IMAGHarmony, a structure-aware framework that incorporates harmony-aware attention (HA) to integrate multimodal semantics, explicitly modeling object counts and layouts to enhance editing accuracy and structural consistency. In addition, we observe that diffusion models are susceptible to initial noise and exhibit strong preferences for specific noise patterns. Motivated by this, we present a preference-guided noise selection (PNS) strategy that chooses semantically aligned initial noise samples based on vision-language matching, thereby improving generation stability and layout consistency in multi-object editing. To support evaluation, we construct HarmonyBench, a comprehensive benchmark covering diverse quantity and layout control scenarios. Extensive experiments demonstrate that IMAGHarmony consistently outperforms state-of-the-art methods in structural alignment and semantic accuracy. The code and model are available at https://github.com/muzishen/IMAGHarmony.
@@ 1 Introduction
Most prompt-driven image editing methods [68, 3, 32, 54, 29, 31] fall into two main categories: mask-free [68, 3, 57, 17, 20] and mask-based approaches [2, 69, 65, 71, 49]. Mask-free approaches [68, 3, 57] rely only on textual prompts to steer the diffusion model, producing global changes without explicit spatial constraints. Mask-based approaches [2, 69, 65], in turn, confine the edit to user-defined masks, enabling fine-grained updates space between text and reference image. As shown in Figure 1 (a), both paradigms perform well when editing a few-object (e.g., one or two). However, they [41, 29, 42, 32] often fail in scenes requiring consistent control over more than two objects, particularly in terms of object counts and spatial layouts from Figure 1 (b). In addition, multi-object editing frequently results in incorrect styles (e.g., Cyberpunk Style) and disorganized backgrounds (e.g., Flowers in Background). Two primary factors contribute to this limitation: (i) The weak counting capability of diffusion models, where text prompts often fail to convey or retain the desired object quantity; (ii) The inability to distinguish multiple entities from object parts within masked regions, hindering accurate structural and semantic modeling.
@@
Furthermore, recent studies [14, 26, 35, 63, 38] have shown that diffusion models are highly susceptible to the choice of initial noise in generation and editing tasks. Since the denoising trajectory [56, 66, 55] is conditioned on the starting noise, different seeds induce distinct attention pathways and layout biases during early stages, leading to substantial variations in object count, spatial arrangement, and visual semantics. For example, InitNO [14] identifies suboptimal noise as a significant cause of prompt-image misalignment and proposes an attention-based strategy to guide noise toward semantically aligned outputs. Building on these insights [64, 63, 36], we observe that initial noise is equally critical in multi-object editing, particularly in determining object count and spatial arrangement. Even under identical prompts, different seeds can produce structurally divergent outputs: some yield clean and well-distributed layouts, while others result in missing, overlapping, or misaligned objects. Therefore, identifying effective noise seeds during inference is crucial for ensuring reliable layout and object consistency in multi-object editing.
@@
To address the above challenges, we present a new task: quantity-and-layout consistent image editing (QL-Edit). Given a reference image and a textual instruction, the objective is to edit the image such that the output not only aligns with the desired semantics but also strictly conforms to the specified object count and spatial arrangement. To this end, we propose IMAGHarmony, a consistency-aware framework that integrates two key components: a harmony-aware attention (HA) module and a preference-guided noise selection (PNS) strategy. Specifically, the HA module addresses the limitations of existing models in quantity and layout modeling by explicitly encoding object count from auxiliary textual features and implicitly capturing spatial relations from visual features. These representations are fused and injected into the denoising network to guide layout-consistent editing. To mitigate the sensitivity of diffusion models to initial noise, the PNS strategy generates multiple candidate seeds during training. At inference, it performs a few-steps denoising for each seed, evaluates semantic alignment using a frozen vision-language model, and selects the best-performing seed to initialize the complete denoising process, thereby improving stability and structural fidelity. To support evaluation, we further construct HarmonyBench, a comprehensive benchmark tailored for QL-Edit. It includes a wide range of multi-object editing scenarios with diverse object counts, spatial layouts, and editing instructions, enabling systematic assessment of both semantic alignment and structural controllability. Our main contributions are summarized as follows:
@@
    We introduce QL-Edit, a new task that enables control image editing while ensuring consistency in object count, category, and spatial layout.
@@
    We propose the IMAGHarmony framework, which integrates a HA module and a PNS strategy to enhance structural consistency and generation stability.
@@
    We construct HarmonyBench, a benchmark covering diverse quantities, categories, and layouts, along with metrics for structural and semantic evaluation.
@@
    Extensive experiments show that IMAGHarmony achieves superior performance over state-of-the-art methods in object count accuracy, layout fidelity, and overall generation quality.
@@ 2 Related Work
Controllable Image Editing with Diffusion Models. Diffusion models [7, 43, 48, 50, 51, 45, 13] have shown strong potential for high-fidelity and semantically aligned image generation and editing. Prompt-driven editing methods can be broadly categorized into two paradigms: (1) mask-free editing [68, 3, 17] and (2) mask-based editing. Mask-free approaches [3, 17, 20] rely solely on natural language instructions to guide image transformation without explicit spatial constraints. For instance, MagicBrush [68] introduces a large-scale manually annotated dataset for real-world instruction-based editing. InstructPix2Pix [3] trains a conditional diffusion model on synthetic image-instruction pairs generated via GPT-3 [10] and Stable Diffusion, enabling efficient instruction-following. Prompt-to-Prompt [17] and SmartEdit [20] modify prompts during inference to achieve semantically aligned edits without requiring additional training. Mask-based approaches [65, 71, 49] provide spatial control by restricting edits to user-defined or automatically generated regions. LEDITS [57] combines DDPM inversion with semantic guidance to support efficient and flexible real-image editing without architectural modifications. LEDITS++[2] further introduces implicit masking and multi-edit support in a tuning-free, architecture-agnostic setting. InstDiffEdit[71] leverages cross-modal attention to enable instant mask generation without training. Click2Mask [49] simplifies localized editing using point-based mask generation and masked CLIP-based semantic loss. Despite these advances, existing methods [3, 17, 20, 65, 71, 49] usually struggle with multi-object editing, failing to control object count and spatial layout due to weak counting ability and limited structural modeling.
@@
Semantic Guidance with Vision-Language Models. Vision-language models (VLMs) [5, 33, 25] have been increasingly employed to enhance semantic controllability in image editing [44, 52, 70, 40, 16]. Early efforts focus on layout-aware guidance, where LayoutLLM-T2I [47] integrates LLM-driven layout planning with object-aware diffusion to achieve high-fidelity image generation without manual spatial annotations. Similarly, LayoutGPT [9] translates textual descriptions into explicit spatial layouts, improving the model’s comprehension of spatial and numerical semantics. Beyond layout guidance, recent studies explore unified visual-linguistic modeling to advance cross-modal understanding. Emu3 [59] constructs a shared token sequence space that supports multi-turn reasoning and progressive image refinement. UniToken [21] and SEED-X [11] introduce transferable token representations that bridge diffusion inversion with structured language decoding, while Show-o [61] improves semantic consistency via a shared multimodal latent space. In parallel, VLMs have also been leveraged at inference time as scoring functions for semantic alignment and content fidelity. InitNO [14] identifies suboptimal seeds by analyzing attention patterns, whereas LS [63] proposes latent space density as an intrinsic indicator of sample quality. PickScore [23] further simulates human preferences via VLM-based ranking to select semantically aligned outputs without human supervision. However, existing methods [47, 15, 60, 61] typically require extensive training overhead and still struggle to align with user intent or produce stable editing results.
Figure 2: Overview of the IMAGHarmony framework. Given a reference image and a text instruction, we first sample multiple candidate noise seeds and evaluate their semantic alignment using a vision-language model (VLM). The top-$k$ candidates are retained for inference. The harmony-aware attention (HA) module fuses auxiliary textual and visual features to jointly model object count and spatial layout. An image prompt (IP) attention layer injects these layout-aware features into the UNet backbone without modifying its frozen weights. During inference, the preference-guided noise selection (PNS) strategy selects the best candidate seed to guide the full denoising process, leading to edited images that are consistent in both quantity and layout.
@@ 3 Method
Quantity-and-layout consistent image editing (QL-Edit) is a novel task that focuses on controllable image manipulation based on textual instructions, while strictly maintaining the original object count and spatial layout from a reference image. Unlike traditional image editing that prioritizes semantic alignment or stylistic fidelity, QL-Edit places equal emphasis on structural integrity, ensuring that modifications do not alter the number or arrangement of key visual elements. To comprehensively evaluate this task, we decompose QL-Edit into the following three representative editing cases:
(1) Class Editing: Alters the object category while preserving the original count and layout (e.g., “Three Apples $\rightarrow$ Three Pineapples”, where both belong to the fruit class, see Figure 1, second row).
(2) Scene Editing: Modifies the background or contextual environment while maintaining object identity, count, and layout (e.g., +“Flowers in Background”, see Figure 1, third row, left).
(3) Style Editing: Applies a stylistic transformation to the image without affecting semantic content or spatial configuration (e.g., +“Cyberpunk Style”, see Figure 1, third row, right).
@@ 3.1 Overall Framework
To address the challenge of controlling object counts and spatial layouts in multi-object image editing, we propose IMAGHarmony, a dedicated framework for the QL-Edit task. As illustrated in Figure 2, IMAGHarmony is built upon the denoising UNet architecture of frozen Stable Diffusion XL (SDXL), and incorporates the proposed harmony-aware attention (HA). Additionally, we present a preference-guided noise selection (PNS) strategy at inference time to enhance structural stability and semantic consistency. Specifically, during training, we first generate multiple initial noise seeds and use SDXL to perform denoising conditioned on the given textual prompt, producing a set of candidate images. These candidates are then evaluated by a frozen vision-language model (VLM) for semantic alignment, and the top-$k$ scoring images along with their corresponding seeds are selected for further processing. The auxiliary text and source image are respectively encoded using frozen text and image encoders to extract conditional text and visual features, which are then fed into the HA module (Section  3.2) to obtain harmony feature that explicitly model object count and implicitly capture layout. Subsequently, a learnable IP attention layer is inserted into the "Down4" block of the UNet to inject the harmony feature. To preserve the generative capability of the original model, we additionally apply cross-attention using a “< Null Prompt >” over the frozen UNet. During inference, we adopt the PNS strategy (Section 3.3) to perform a few denoising steps on the candidate seeds, evaluate their semantic consistency, and ultimately select the optimal noise seed to conduct full denoising, resulting in high-quality edited images that align with the desired object count and spatial layout. Finally, we provide a comprehensive summary of the training and inference strategies implemented in IMAGHarmony (Section 3.4).
3.2 Harmony-Aware Attention Module
Motivation. Diffusion models often struggle to accurately capture object count and spatial layout from visual features alone, due to the absence of explicit modeling for quantity and spatial relations. This limitation hinders their effectiveness in structured multi-object editing. To address this, we propose a harmony-aware attention (HA) module that integrates auxiliary textual and visual cues to encode object count and implicitly capture layout. The HA module facilitates more consistent and controllable editing outcomes by enhancing the model’s perception of quantity and structure.
Architecture. As illustrated on the right of Figure 2 (a), the proposed HA module consists of two learnable linear projections and a multi-head cross-attention mechanism, designed to enhance cross-modal alignment of semantic and structural information. Specifically, let $F_{v}\in\mathbb{R}^{B\times N\times D_{v}}$ and $F_{a}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the visual and auxiliary textual features, where $B$ is the batch size, $N$ the number of tokens. $D_{v}$ and $D_{t}$ are the respective feature dimensions. We first project visual feature $F_{v}$ into the auxiliary textual feature space, according to Eq. (1).

### v2

@@ Abstract
Recent diffusion models have advanced image editing by improving fidelity and controllability across creative and personalized applications. However, multi-object scenes remain challenging, as reliable control over object categories, counts, and spatial layout is difficult to achieve. For that, we first study quantity and layout consistent image editing, abbreviated as QL-Edit, which targets control of object quantity and spatial layout in multi-object scenes. Then, we present IMAGHarmony, a straightforward framework featuring a plug-and-play harmony aware (HA) module that fuses perception semantics while modeling object counts and locations, resulting in accurate edits and strong structural consistency. We further observe that diffusion models are sensitive to the choice of initial noise and tend to prefer certain noise patterns. Based on this finding, we present a preference-guided noise selection (PNS) strategy that selects semantically aligned initial noise through vision and language matching, thereby further improving generation stability and layout consistency in multiple object editing. To support evaluation, we develop HarmonyBench, a comprehensive benchmark that covers a diverse range of quantity and layout control scenarios. Extensive experiments demonstrate that IMAGHarmony outperforms prior methods in both structural alignment and semantic accuracy, utilizing only 200 training images and 10.6M of trainable parameters. Code, models, and data are available at https://github.com/muzishen/IMAGHarmony.
Index Terms:
Image Editing, Diffusion Models, Controllable Generation, Object Counting, Layout Consistency.
@@ 1 Introduction
Most prompt-driven image editing methods [18, 19, 17, 20, 21, 22] fall into two main categories: mask-free [18, 19, 23, 24, 2] and mask-based approaches [25, 26, 27, 28, 29]. Mask-free approaches [18, 19, 23] rely only on textual prompts to steer the diffusion model, producing global changes without explicit spatial constraints. Mask-based approaches [25, 26, 27], in turn, confine the edit to user-defined masks, enabling fine-grained updates space between text and reference image. As shown in Fig. 1 (a), both paradigms perform well when editing a few-object (e.g., one or two). However, they [15, 21, 16, 17] often fail in scenes requiring consistent control over more than two objects, particularly in terms of object counts and spatial layouts from Figure 1 (b). In addition, multi-object editing frequently results in incorrect styles (e.g., + Cyberpunk Style) and disorganized backgrounds (e.g., in Flowers). This is primarily due to the weak counting ability of diffusion models: text prompts often fail to convey or preserve the required object count, and the model struggles to disambiguate multiple entities and object parts within masked regions, which in turn hinders accurate structural and semantic modeling.
@@
Furthermore, recent studies [30, 31, 3, 32, 33] show that diffusion models are highly sensitive to the choice of the initial noise during inference, in both generation and editing tasks. Since the denoising trajectory [34, 35, 36] is conditioned on the starting noise, different seeds induce distinct attention pathways and layout biases during early stages, leading to substantial variations in object count, spatial arrangement, and visual semantics. For example, InitNO [30] identifies suboptimal noise as a significant cause of prompt-image misalignment and proposes an attention-based strategy to guide noise toward semantically aligned outputs. In contrast, LS [32] highlights the limitations of using external feature extractors for quality evaluation and instead introduces latent space density as a model-intrinsic indicator of generation quality. Building on these insights [37, 32, 38], we observe during training that the choice of the initial noise seed is equally critical for multi-object editing, particularly in governing object count and spatial arrangement. Even under identical prompts, different seeds can produce structurally divergent outputs: some yield clean and well-distributed layouts, while others result in missing, overlapping, or misaligned objects. Therefore, identifying effective noise seeds during training and inference is crucial for ensuring reliable layout and object consistency in multi-object editing. To sum up, there are mainly two challenges: (i) weak counting and part–instance ambiguity, and (ii) high sensitivity to initial noise that destabilizes trajectories.
@@
For that, we first study quantity and layout consistent image editing, abbreviated as QL-Edit, where a reference image and a textual instruction are used to produce an edited image that aligns with the target semantics while strictly matching the specified object count and spatial arrangement. Then, we propose IMAGHarmony, a consistency-aware framework that combines a harmony aware (HA) module with a preference-guided noise selection (PNS) strategy. The HA module consists of a quantity-and-layout (QL) attention and a cross attention layer. The former explicitly encodes object counts from auxiliary textual cues and implicitly captures spatial relations from visual features; it then performs cross-modal interaction and complementarity to produce the perception features. These perception features are subsequently fused via a cross attention and injected into the denoising network, thereby guiding consistent editing. To reduce sensitivity to the initial noise, the preference-guided noise selection (PNS) strategy generates multiple candidate seeds during training and inference, performs noising and denoising for each, evaluates semantic consistency with a frozen vision–language model (VLM), and selects the best-performing seed to initialize the full diffusion trajectory, thereby improving stability and structural fidelity. For evaluation, we construct HarmonyBench, a comprehensive benchmark that covers diverse multiple object scenarios with varied counts, spatial layouts, and editing instructions, enabling systematic assessment of semantic alignment and structural controllability. Extensive experiments show that IMAGHarmony surpasses state-of-the-art methods in structural alignment and semantic accuracy using only 200 training images and just 10.6M trainable parameters.
The main contributions of this paper are summarized as follows:
@@
    We present IMAGHarmony for controllable image editing in QL-Edit, ensuring consistent object counts, categories, and spatial layouts in multi-object scenes.
@@
    We design a HA module that explicitly encodes object counts and implicitly captures spatial relations, injecting these cues into the denoising network to enable consistent editing.
@@
    We devise a PNS strategy that, during training and inference, selects semantically aligned noise seeds via vision and language matching to stabilize the diffusion trajectory and preserve structural fidelity.
@@
    We construct HarmonyBench, a multi-object editing benchmark, and show that IMAGHarmony achieves state-of-the-art performance in QL-Edit.
The remainder of this paper is organized as follows. Section 2 reviews related work on image editing and VLM. Section 3 presents the proposed IMAGHarmony, including preliminaries, the overall framework, and its two core components: the HA module and the PNS strategy. Section 4 reports experiments and analyses on standard benchmarks. Finally, Section 5 concludes this work.
@@ 2 Related Work
2.1 Controllable Image Editing with Diffusion Models
@@
Diffusion models [7, 39, 8, 9, 40, 41, 42] have shown strong potential for high-fidelity and semantically aligned image generation and editing. Prompt-driven editing methods can be broadly categorized into two paradigms: (1) mask-free editing [18, 19, 24] and (2) mask-based editing. Mask-free approaches [19, 24, 2] rely solely on natural language instructions to guide image transformation without explicit spatial constraints. For instance, MagicBrush [18] introduces a large-scale manually annotated dataset for real-world instruction-based editing. InstructPix2Pix [19] trains a conditional diffusion model on synthetic image-instruction pairs generated via GPT-3 [43] and Stable Diffusion, enabling efficient instruction-following. Prompt-to-Prompt [24] and SmartEdit [2] modify prompts during inference to achieve semantically aligned edits without requiring additional training. Mask-based approaches [27, 28, 29] provide spatial control by restricting edits to user-defined or automatically generated regions. Furthermore, LEDITS [23] combines DDPM inversion with semantic guidance to support efficient and flexible real-image editing without architectural modifications. LEDITS++[25] further introduces implicit masking and multi-edit support in a tuning-free, architecture-agnostic setting. Then, InstDiffEdit[28] leverages cross-modal attention to enable instant mask generation without training. Click2Mask [29] simplifies localized editing using point-based mask generation and masked CLIP-based semantic loss.
Recent works [44, 1, 45] have likewise shown that diffusion models exhibit limited controllability over multiple objects during generation. For example, YOLO-Count [44] introduces an auxiliary YOLO detector to bridge the gap between count control and text-to-image (T2I) generation. Despite such advances, in the editing settingmost existing methods [19, 24, 2, 27, 28, 29] still struggle with multi-object cases, often failing to control object counts and spatial layout due to weak counting ability and limited structural modeling. To this end, we focus on and investigate object-count control in diffusion models for image editing.
2.2 Semantic Guidance with Vision-Language Models
@@
Vision-language models (VLMs) [46, 47, 48] have been increasingly employed to enhance semantic controllability in image editing [49, 50, 51, 52, 53]. Early efforts focus on layout-aware guidance, where LayoutLLM-T2I [54] integrates LLM-driven layout planning with object-aware diffusion to achieve high-fidelity image generation without manual spatial annotations. Similarly, LayoutGPT [55] translates textual descriptions into explicit spatial layouts, improving the model’s comprehension of spatial and numerical semantics. Beyond layout guidance, recent studies explore unified visual-linguistic modeling to advance cross-modal understanding. Emu3 [56] constructs a shared token sequence space that supports multi-turn reasoning and progressive image refinement. Furthermore, UniToken [57] and SEED-X [58] introduce transferable token representations that bridge diffusion inversion with structured language decoding, while Show-o [59] improves semantic consistency via a shared multimodal latent space.
In parallel, VLMs have also been leveraged at inference time as scoring functions for semantic alignment and content fidelity. For example, InitNO [30] identifies suboptimal seeds by analyzing attention patterns, whereas LS [32] proposes latent space density as an intrinsic indicator of sample quality. PickScore [60] further simulates human preferences via VLM-based ranking to select semantically aligned outputs without human supervision. However, in unified VLM-based understanding and generation frameworks, existing methods [54, 61, 62, 59] typically rely on explicit layout annotations (e.g., bounding boxes) and demand substantial training data and computing resources. For VLM-driven noise selection, prior work [30, 60] primarily focuses on inference-time seed choice while overlooking the impact of training-time noise on model behavior. Therefore, we propose IMAGHarmony, which eliminates the need for explicit layout labels by capturing layout implicitly from visual features and stabilizes diffusion trajectories by addressing noise at both training and inference stages.
TABLE I: Comparison of other tasks under image editing. QL-Edit targets multi-object, class, scene, and style editing.
Task Multi-Object Class Scene Style
Prompt-to-Prompt
Object-Centric Editing
Layout-to-Image
QL-Edit
@@ 3 Method
In Section 3.1, we formalize the QL-Edit problem, delineate its differences from prior image editing tasks, and introduce the notation used throughout the paper. We next introduce IMAGHarmony, a framework tailored to QL Edit. Section 3.2 and Fig. 2 present the overall pipeline. IMAGHarmony augments a frozen Stable diffusion XL (SDXL) denoising UNet with a harmony aware (HA) module, detailed in Section 3.3. Then, Section 3.4 describes a preference-guided noise selection (PNS) strategy applied during training and inference to improve structural stability and semantic consistency. Section 3.5 summarizes the complete training and inference procedure.
TABLE II: Definitions of main symbols used in this paper.
Notation Definition
$\boldsymbol{x_{0}}$ Target image
$\boldsymbol{\epsilon}$ Gaussian noise
$\boldsymbol{t}$ Timestep
$\boldsymbol{T}$ Total number of denoising steps
$\boldsymbol{\theta}$ Diffusion model
$\boldsymbol{w}$ Guidance scale
$\boldsymbol{\alpha}$ Scaling hyperparameter of harmony attention
$\boldsymbol{\beta}$ IP attention scale
$\boldsymbol{x_{t}}$ Noisy data at $t$ step
$\boldsymbol{Z_{t}}$ Noisy latent
$\boldsymbol{F_{v}}$ Visual feature
$\boldsymbol{F_{a}}$ Auxiliary textual feature
$\boldsymbol{F_{t}}$ Textual feature
$\boldsymbol{F_{p}}$ Perception feature
$\boldsymbol{S_{top}}$ Candidate seed set
$\boldsymbol{\hat{I}_{j}}$ Candidate image
$\boldsymbol{A_{i}}$ Binary accuracy score
$\boldsymbol{R_{j}}$ Semantic alignment score
Figure 2: Overview of the IMAGHarmony framework. Given a source image and a text instruction, we first sample multiple candidate noise seeds and judge their semantic alignment using a vision-language model (VLM). The top-$k$ candidates are selected for inference. The harmony aware (HA) module fuses auxiliary textual and visual features to jointly model object count and spatial layout. Then, a cross attention layer injects these perception features into the UNet backbone without modifying its frozen weights. The preference-guided noise selection (PNS) strategy searches for and selects the best candidate seed to stabilize trajectories across training and inference, ultimately leading to consistent edited images.
3.1 Preliminaries
Problem Description. We study quantity and layout consistent image editing, abbreviated as QL-Edit, which performs controllable manipulation from textual instructions while strictly preserving the object count and spatial arrangement in a reference image. Unlike conventional editing that primarily targets semantic alignment or style, QL-Edit gives equal weight to structural integrity so that edits do not change the number or the placement of key visual elements. We decompose QL-Edit into three representative cases, as follows: (1) Class Editing: Alters the object category while preserving the original count and layout (e.g., “Six Tigers $\rightarrow$ Six Lions”, where both belong to the animal class, see Fig. 1, second row). (2) Scene Editing: Modifies the background or contextual environment while maintaining object count, and layout (e.g., +“Flowers in Background”, see Fig. 1, third row, left). (3) Style Editing: Applies a stylistic transformation to the image without affecting semantic content or spatial configuration (e.g., +“Cyberpunk Style”, see Fig. 1, third row, right). Therefore, QL-Edit denotes image editing in multi-object scenes under the constraint that the edited result preserves the source image’s quantity–layout consistency, rather than simply executing generic instruction-based edits.
Task Differences. Multi-object image editing is pervasive in real-world applications, where scenes often contain three or more objects. Yet, as summarized in TABLE I, existing directions either provide limited editing coverage or lack structural constraints. Prompt-to-Prompt [18] flexibly handles class and style transformations but neither models multi-object behavior nor enforces layout preservation. Object-centric editing [19] typically operates on local regions and overlooks scene-level coherence. Layout-to-image [55] relies on predefined layouts and omits quantity control. In contrast, QL-Edit imposes joint constraints on quantity and layout while supporting class, scene, and style editing.
Symbol Definition. To introduce our proposed IMAGHarmony method more clearly, we define the main symbols used throughout the paper in TABLE II.
@@ 3.2 Overall Framework
As shown in Fig. 2, we first generate multiple initial noise seeds and use SDXL to perform denoising conditioned on the given textual prompt, producing a set of candidate images. These candidates are then evaluated by a frozen vision-language model (VLM) for semantic alignment, and the top-$k$ scoring images along with their corresponding seeds are selected for further processing. The auxiliary text and source image are respectively encoded using frozen text and image encoders to extract conditional text and visual features, which are then fed into the HA module to obtain harmony feature that explicitly model object count and implicitly capture layout. Subsequently, a learnable cross attention layer is inserted into the "Down4" block of the UNet to inject the harmony feature. To preserve the generative capability of the original model, we additionally apply frozen cross attention using a “< Null Prompt >” over the frozen UNet. During inference, we adopt the PNS strategy to perform a few denoising steps on the candidate seeds, evaluate their semantic consistency, and ultimately select the optimal noise seed to conduct full denoising, resulting in high-quality edited images that align with the desired object count and spatial layout. Finally, we provide a comprehensive summary of the training and inference strategies implemented in IMAGHarmony.
3.3 Harmony Aware Module
Motivation. Existing work [44, 63] indicates that diffusion models often struggle to accurately capture object count and spatial layout from visual features alone, due to the absence of explicit modeling for quantity and spatial relations. This limitation hinders their effectiveness in structured multi-object editing. To address this, we propose a harmony aware (HA) module that integrates auxiliary textual and visual cues to encode object count and implicitly capture layout. The HA module facilitates more consistent and controllable editing outcomes by enhancing the model’s perception of quantity and layout.
Architecture. As illustrated in Fig. 2 (a), the proposed HA module comprises a quantity-and-layout (QL) attention and a cross attention layer. The QL attention contains two learnable linear projections and a multi-head cross attention mechanism, designed to enhance cross-modal perception of semantic and structural information. Specifically, let $F_{v}\in\mathbb{R}^{B\times N\times D_{v}}$ and $F_{a}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the visual and auxiliary textual features, where $B$ is the batch size, $N$ the number of tokens. $D_{v}$ and $D_{t}$ are the respective feature dimensions. We first project visual feature $F_{v}$ into the auxiliary textual feature space, according to Eq. (1).